OpenStudy (anonymous):

A certain nasty bacterium splits every 1.2 seconds (into two nasty bacteria). These bacteria never die. a) What is the doubling time of a population of these bacteria? b) If there are currently 4294967296 bacteria in a population, how long ago was there just a single bacterium? I'm having trouble with the second problem, it must be inverse of 2^n = 1.2n but how would i do that

OpenStudy (amistre64):

when does 2^n = that big number?

OpenStudy (amistre64):

log both sides and grease up the gears in your calculator :)

OpenStudy (amistre64):

2^n = 4294967296 L(2^n) = L(4294967296) n L(2) = L(4294967296) n = L(4294967296)/L(2)

OpenStudy (anonymous):

o wow i didn't even think to use logs

OpenStudy (anonymous):

saddest part is that i just did a problem that involved logs T___T thanks amistre64!

OpenStudy (amistre64):

with any luck, you get n=32 ;)